The manager of a discount clothing store received two shipments of fall clothing. The cost of the first shipment which contained ten identical sweaters and twenty identical jackets was $800. The second shipment at the same prices, contained five of the same sweaters and fifteen of the same jackets. The second shipment was $550. Find the cost of one jacket.
Answer by checkley77(12569)

20J+10S=800
15J+5S=550 MULTIPLY THIS EQUATION BY -2 & ADD.
-30J-10S=-1100
20J+10S=800
---------------------
-10J=-300
J=-300/-10
J=$30 THE PRICE OF EACH JACKET.
20*3+10X=800
600+10X=800
10X=800-600
10X=200
X=200/10
X=$20 FOR EACH SWEATER.
PROOF:
15*30+5*20=550
450+100=550